# An interpretable family of projected normal distributions and a related copula model for Bayesian analysis of hypertoroidal data

### v1

@@ 3.2 An extension of the copula case
Let a random vector $(U_{1},\ldots,U_{d})$ follow the proposed model $TPN_{d}(\bm{0},\bm{0},\bm{\Sigma})$. Let $f_{m,j}(\theta),\ j=1,\ldots,d,$ be densities on the circle with the mode at $\theta=0$ and $F_{m,j}(\theta)=\int_{0}^{\theta}f_{m,j}(t)dt$ be their cumulative distribution functions (CDFs). Define the circular random variables by
$\Theta_{1}=\mu_{1}+F_{m,1}^{-1}\left(\frac{U_{1}}{2\pi}\right),\quad\ldots,\quad\Theta_{d}=\mu_{d}+F_{m,d}^{-1}\left(\frac{U_{d}}{2\pi}\right).$ (19)
@@
It immediately follows from Sklar’s theorem (e.g., Nelsen,, 2006, Theorem 2.10.9) that the random vector $(\Theta_{1},\ldots,\Theta_{d})$ defined in Definition 2 has the probability density function
$f(\bm{\theta})=(2\pi)^{d}\overline{c}(2\pi F_{m,1}(\theta_{1}-\mu_{1}),\ldots,2\pi F_{m,d}(\theta_{d}-\mu_{d}))\prod_{j=1}^{d}f_{m,j}(\theta_{j}-\mu_{j}),$ (20)
$\bm{\theta}=(\theta_{1},\ldots,\theta_{d})^{\top}\in[-\pi,\pi)^{d},$
@@
$f(\theta_{j})=f_{m,j}(\theta_{j}-\mu_{j}).$
The assumption that the mode of $f_{m,j}(\theta)$ is at $\theta=0$ ensures that the mode of each marginal of the CTPN distribution is given by $\theta_{j}=\mu_{j}$. The lower limit of the CDF $F_{m,j}$ is assumed to be zero because the mode of the joint density $f(\bm{\theta})$ of the CTPN distribution is located at $\bm{\theta}=(\mu_{1},\ldots,\mu_{d})$ if the copula density $\overline{c}(\bm{u})$ has the mode at $\bm{u}=\bm{0}$ (see Jones et al.,, 2015, Section 3.1). Flexible modeling is made possible through this extended model (20).
@@
$f(\theta_{j},\theta_{k})=(2\pi)^{2}\overline{c}_{2}(2\pi F_{m,j}(\theta_{j}),2\pi F_{m,k}(\theta_{k}))f_{m,j}(\theta_{j})f_{m,k}(\theta_{k}),$ (21)
where $\overline{c}_{2}$ denotes the density (18) with $\mu_{j}=\mu_{k}=0$. This bivariate case belongs to the general family of Wehrly and Johnson, (1980) with the circular density binding the two variables being $2\pi\overline{c}_{2}(\theta_{j},\theta_{k})$.
@@ 4 Coefficient of Correlation
Figure 3: TNP distribution: plot of the circular correlation for different values of $\rho_{jk}$, and values of $\kappa_{j},\kappa_{k}\in\{0,0.49,2.45\}$: (a) $\kappa_{j}=\kappa_{k}$; and (b) $\kappa_{j}\neq\kappa_{k}$.
As seen in Figure 2, the parameter $\rho_{jk}$ governs the dependence between $\Theta_{j}$ and $\Theta_{k}$. To quantify how the value of $\rho_{jk}$ influences the dependence, we consider the circular correlation coefficient proposed by Rivest, (1982). This measure is considered here because it is a signed coefficient capable of distinguishing between positive and negative dependence. Let
$\xi_{jk}=\frac{\left|E(e^{i(\Theta_{j}-\Theta_{k})})\right|-\left|E(e^{i(\Theta_{j}+\Theta_{k})})\right|}{2},$
then the circular correlation coefficient between $\Theta_{j}$ and $\Theta_{k}$ proposed by Rivest, (1982) is defined as
$\rho_{jk,R}=\left\{\begin{array}[]{ll}\frac{\xi_{jk}}{\max\left\{E\left[\sin^{2}(\Theta_{j}-\mu_{j})\right],\ E\left[\sin^{2}(\Theta_{k}-\mu_{k})\right]\right\}},&E(e^{i(\Theta_{j}-\Theta_{k})})E(e^{i(\Theta_{j}+\Theta_{k})})\neq 0,\\ 2\xi_{jk},&E(e^{i(\Theta_{j}-\Theta_{k})})E(e^{i(\Theta_{j}+\Theta_{k})})=0.\end{array}\right$.
Figure 3 shows the circular correlation coefficient $\rho_{jk,R}$ for two cases: $\kappa_{j}=\kappa_{k}$ and $\kappa_{j}\neq\kappa_{k}$. The figure suggests that the circular correlation coefficient is a strictly increasing function of $\rho_{jk}$. The coefficient spans the entire range (1,1)(-1,1) when $\kappa_{j}=\kappa_{k}$, whereas its range is more limited when $\kappa_{j}\neq\kappa_{k}$. This can be inferred from the fact that if the range of $\rho_{jk}$ is extended and $\rho_{jk}=-1$ or $1$ in Definition 1, we have
@@
$\Theta_{k}=\mu_{k}+\operatorname{sign}(\rho_{jk})\,\text{atan}^{*}\left(\frac{\sin(\Theta_{j}-\mu_{j})}{\cos(\Theta_{j}-\mu_{j})+(\kappa_{k}-\kappa_{j})/R_{j}}\right).$
This expression implies that $\Theta_{j}$ and $\Theta_{k}$ have perfect linear relationship when $\kappa_{j}=\kappa_{k}$. When $\kappa_{j}\neq\kappa_{k}$, the relationship between $\Theta_{j}$ and $\Theta_{k}$ depends on $R_{j}$, and the influence of $R_{j}$ increases as $|\kappa_{k}-\kappa_{j}|$ increases. This suggests the advantage of the CTPN distribution over the TPN distribution. Note, however, that the circular correlation coefficient $\rho_{jk,R}$ reaches 1-1 or $1$ when $\Theta_{j}$ and $\Theta_{k}$ have a linear relationship with probability one (see Rivest,, 1982, Proposition 2.3). Therefore, this coefficient is not suitable for measuring the relationship between the two variables when they exhibit a nonlinear dependence, which may occur when their marginal distributions differ.
In addition, it can be readily confirmed from Figure 3 that the positive and negative signs of $\rho_{jk}$ correspond to positive and negative values of the circular correlation coefficient, respectively.
@@ 5 MCMC Algorithm
In this section, we explain the algorithm that can be used to obtain posterior samples from the TPN distribution. To deal with the identification problem that the TPN inherited form the PN, we use the approach of Mastrantonio, (2018), which defines a model for the unidentifiable parameters and a posterior remap is used to obtain the identifiable version of the parameters.
While standard priors can be easily assigned to the parameters $\bm{\mu}$ and $\bm{\kappa}$, care must be taken when specifying a prior for $\bm{\Sigma}$ due to its relationship with $\bm{\Sigma}_{c}$. The Inverse Wishart (IW) distribution is a classical choice for modeling a covariance matrix, but not every valid covariance matrix $\bm{\Sigma}$ yields a valid $\bm{\Sigma}_{c}$. This makes its use problematic unless appropriate adjustments are made. The matrix $\bm{\Sigma}$ is defined over the domain $\mathcal{D}$ of all positive definite matrices, and a subset $\mathcal{D}^{*}$ of this domain can be defined as the set of all positive matrices that, when taking the absolute value of their elements, remain positive definite. We define a new distribution, which we call Truncated IW (TIW), as a distribution with value over $\mathcal{D}^{*}$ that has a density that is proportional to the IW. The normalization constant of this distribution is not known, but, as we will show, it is not needed in the MCMC algorithm.
@@
$\bm{\Theta}_{i}$ $\stackrel{{\scriptstyle iid}}{{\sim}}TPN_{d}(\bm{\mu},\bm{\kappa},\bm{\Sigma}),$ (22)
$\bm{\Sigma}$ $\sim TIW(\nu,\bm{\Psi}),\quad i=1,\dots,n,$ (23)
$\mu_{j}$ $\stackrel{{\scriptstyle iid}}{{\sim}}CU(0,2\pi),$ (24)
$\kappa_{j}$ $\stackrel{{\scriptstyle iid}}{{\sim}}N(\mathbf{m},\mathbf{v})I_{>0},\quad j=1,\dots,d,$ (25)
where $N(\mathbf{m},\mathbf{v})I_{>0}$ denotes a normal distribution with mean $\mathbf{m}$ and variance $\mathbf{v}$ truncated to the positive real line, while $CU(0,2\pi)$ is the uniform distribution on the circle $(0,2\pi)$. As we will show below, the truncated normal ensures that the full conditional of $\kappa_{j}$ is available in closed form, while for $\mu_{j}$, regardless of the distribution, no closed-form full conditional is available.
To make the inference easier, as is often done when working with the PN, the vectors of latent variables $\mathbf{R}_{i}=(R_{i,1},\dots,R_{i,d})^{\top}$ are introduced in the model. Then by defining the $j$-th element of $\mathbf{X}_{c}$ and $\mathbf{X}_{s}$ as
$[\mathbf{X}_{c}]_{j}$ $=R_{i,j}\cos(\Theta_{i,j}-\mu_{j}),\quad[\mathbf{X}_{s}]_{j}=R_{i,j}\sin(\Theta_{i,j}-\mu_{j}),$ (26)
we have that the joint density of $(\mathbf{R}_{i},\bm{\Theta}_{i})$ can be written as in (10), as
@@
5.1 MCMC steps for the TPN
For two of the steps of the MCMC algorithm, we need to use the conditional distribution of $x_{c,i,j}$ given all the other variables in the vector $\bm{x}_{c,i}$, which we indicate as $\bm{x}_{c,i,-j}$, and the conditional distribution of $x_{s,i,j}|\bm{x}_{s,i,-j}$. If we define $\bm{\Lambda}_{c}=\bm{\Sigma}_{c}^{-1}$, $\bm{\Lambda}_{s}=\bm{\Sigma}^{-1}$ and $\bm{\kappa}_{-j}$ as the vector $\bm{\kappa}$ without the $j$-the element, using standard results of a block matrix inversion and its connection with the conditional mean and variance of a multivariate normal, we know that
$x_{c,i,j}|\bm{x}_{c,i,-j}$ $\sim N(\eta_{c,i,j},\lambda_{c,j}^{-1}),\quad x_{s,i,j}|\bm{x}_{s,i,-j}\sim N(\eta_{s,i,j},\lambda_{s,j}^{-1}),$ (28)
$\eta_{c,i,j}$ $=\kappa_{j}-\lambda_{c,j}^{-1}\bm{\Lambda}_{c,j,-j}\left(\bm{x}_{c,i,-j}-\bm{\kappa}_{-j}\right),\quad\eta_{s,i,j}=-\lambda_{s,j}^{-1}\bm{\Lambda}_{s,j,-j}\bm{x}_{s,i,-j}$ (29)
@@ 5.1 MCMC steps for the TPN
Latent variable $r_{i,j}$.
We can easily see that the full conditional of $r_{i,j}$ is proportional to
$\exp\left[-\frac{(x_{c,i,j}-\eta_{c,i,j})^{2}}{2\lambda_{c,j}^{-1}}\right]\exp\left[-\frac{(x_{s,i,j}-\eta_{s,i,j})^{2}}{2\lambda_{s,j}^{-1}}\right]r_{i,j}$ (30)
which, by defining $\bm{\eta}_{i,j}=(\eta_{c,i,j},\eta_{s,i,j})^{\top}$, $\tilde{\bm{\Sigma}}_{i,j}=\text{diag}(\lambda_{c,j}^{-1},\lambda_{s,j}^{-1})^{\top}$, and $\mathbf{x}_{i,j}=(x_{c,i,j},x_{s,i,j})^{\top}$, can be written as
$\exp\left[-\frac{(\mathbf{x}_{i,j}-\bm{\eta}_{i,j})^{\top}\tilde{\bm{\Sigma}}_{i,j}^{-1}(\mathbf{x}_{i,j}-\bm{\eta}_{i,j})}{2}\right]r_{i,j}$ (31)
which shows that $\bm{\theta}_{i,j}|\dots\sim PN(\bm{\eta}_{i,j},\tilde{\bm{\Sigma}}_{i,j})$, and hence we can use the algorithm proposed by Hernandez-Stumpfhauser et al., (2017). First, we define
$\bm{\xi}_{i,j}=(\cos(\theta_{i,j}),\sin(\theta_{i,j}))^{\top},\quad A_{i,j}=\mathbf{u}_{i}^{\top}\tilde{\bm{\Sigma}}_{i,j}^{-1}\bm{\xi}_{i,j},\quad B_{i,j}=\bm{\xi}_{i,j}^{\top}\tilde{\bm{\Sigma}}_{i,j}^{-1}\bm{\eta}_{i,j}.$
Then, we draw ${\beta}_{1,i,j}\sim U\left(0,\exp\left(-\frac{1}{2}A_{i,j}\left(r_{i,j}-\frac{B_{i,j}}{A_{i,j}}\right)^{2}\right)\right)$ and ${\beta}_{2,i,j}\sim U(0,1)$, and the new sample of $r_{i,j}$ is
$r_{i,j}=\left((\varrho_{2,i,j}^{2}-\varrho_{1,i,j}^{2}){\beta}_{2,i,j}+\varrho_{1,i,j}^{2}\right)^{1/2},$
@@
$\varrho_{1,i,j}=\frac{B_{i,j}}{A_{i,j}}+\max\left\{-\frac{B_{i,j}}{A_{i,j}},-\sqrt{\frac{-2\log{\beta}_{1,i,j}}{A_{i,j}}}\right\},\quad\varrho_{2,i,j}=\frac{B_{i,j}}{A_{i,j}}+\sqrt{\frac{-2\log{\beta}_{1,i,j}}{A_{i,j}}}.$
Parameter $\mu_{j}$.
Being $\mu_{j}$ embedded inside the sine and cosine used to compute $\mathbf{X}_{c}$ and $\mathbf{X}_{s}$, it is not generally possible to find a closed form expression for its full conditional, which must be then sampled using a Metropolis Hastings algorithm.
Parameter $\kappa_{j}$.
Since the kernel of the prior distribution of $\kappa_{j}$ is the same as the one of a normal distribution, from the conditional distribution (28) we can see that
$x_{c,i,j}+\lambda_{c,j}^{-1}\bm{\Lambda}_{c,j,-j}\left(\bm{x}_{c,i,-j}-\bm{\kappa}_{-j}\right)\sim N(\kappa_{j},\lambda_{c,j}^{-1}).$
Hence, the full conditional of $\kappa_{j}$ is proportional to
$\exp\left[-\frac{1}{2}\left(\kappa_{j}^{2}v^{-1}-2\kappa_{j}v^{-1}m+\kappa_{j}^{2}n\lambda_{c,j}-2\kappa_{j}\sum_{i=1}^{n}\left(\bm{x}_{c,i}+\lambda_{c,j}^{-1}\bm{\Lambda}_{c,j,-j}\left(\bm{x}_{c,i,-j}-\bm{\kappa}_{-j}\right)\right)\right)\right]$ (32)
$\propto\exp\left[-\frac{(\kappa_{j}-m_{p,j})^{2}}{2v_{p,j}}\right],$ (33)
where
$v_{p,j}$ $=(v^{-1}+n\lambda_{c,j})^{-1},$
$m_{p,j}$ $=v_{p,j}\left(\sum_{i=1}^{n}\left(\bm{x}_{c,i}+\lambda_{c,j}^{-1}\bm{\Lambda}_{c,j,-j}(\bm{x}_{c,i,-j}-\bm{\kappa}_{-j})\right)+v^{-1}m\right).$
It is easy to see that this is the kernel of $N(m_{p,j},v_{p,j})$, and since $\kappa_{j}>0$, the full conditional of $\kappa_{j}$ is $N(m_{p,j},v_{p,j})\mathbf{I}_{\kappa_{j}>0}$.
Parameter $\bm{\Sigma}$.
To sample this parameter, we use a Metropolis–Hastings step. At iteration $b$, the proposal distribution is set to $IW(\tau,\bm{\Sigma}^{b}(\tau-(d+1)))$, where $\bm{\Sigma}^{b}$ denotes the most recently accepted value of the parameter in the MCMC. The expected value of this proposal distribution is $\bm{\Sigma}^{b}$, and as the parameter $\tau$ increases, the variance of the distribution decreases, making $\tau$ a tuning parameter that controls the variability of the proposal.

### v2

@@ 3.2 An extension of the copula case
Let a random vector $(U_{1},\ldots,U_{d})$ follow the proposed model $\mathrm{TPN}_{d}(\bm{0},\bm{0},\bm{\Sigma})$. Let $f_{j}(\theta)\ (j=1,\ldots,d)$ be densities on the circle with the mode at $\theta=0$ and $F_{j}(\theta)=\int_{0}^{\theta}f_{j}(t)dt$ be their cumulative distribution functions. Define the circular random variables by
$\Theta_{1}=\mu_{1}+F_{1}^{-1}\left(\frac{U_{1}}{2\pi}\right),\quad\ldots,\quad\Theta_{d}=\mu_{d}+F_{d}^{-1}\left(\frac{U_{d}}{2\pi}\right).$ (17)
@@
It immediately follows from Sklar’s theorem (e.g., Nelsen2006, Theorem 2.10.9) that the random vector $\bm{\Theta}=(\Theta_{1},\ldots,\Theta_{d})^{\top}$ defined in Definition 2 has the probability density function
$f(\bm{\theta})=(2\pi)^{d}\overline{c}\{2\pi F_{1}(\theta_{1}-\mu_{1}),\ldots,2\pi F_{d}(\theta_{d}-\mu_{d})\}\prod_{j=1}^{d}f_{j}(\theta_{j}-\mu_{j})$ (18)
$(\bm{\theta}=(\theta_{1},\ldots,\theta_{d})^{\top}\in[-\pi,\pi)^{d}),$
@@
$f(\theta_{j})=f_{j}(\theta_{j}-\mu_{j}).$
The assumption that the mode of $f_{j}(\theta)$ is at $\theta=0$ ensures that the mode of each marginal of the CTPN distribution is given by $\theta_{j}=\mu_{j}$. The lower limit of $F_{j}$ is assumed to be zero because the mode of the joint density $f(\bm{\theta})$ of the CTPN distribution is located at $\bm{\theta}=(\mu_{1},\ldots,\mu_{d})$ if the copula density $\overline{c}(\bm{u})$ has the mode at $\bm{u}=\bm{0}$; see Jones2015. Flexible modeling is made possible through this extended model (18).
@@
$f(\theta_{j},\theta_{k})=(2\pi)^{2}\overline{c}_{2}\{2\pi F_{j}(\theta_{j}),2\pi F_{k}(\theta_{k})\}f_{j}(\theta_{j})f_{k}(\theta_{k}),$ (19)
where $\overline{c}_{2}$ denotes the density (16) with $\mu_{j}=\mu_{k}=0$. This bivariate case belongs to the general family of Wehrly1980 with the circular density binding the two variables being $2\pi\overline{c}_{2}(\theta_{j},\theta_{k})$.
@@ 4 Coefficient of Correlation
As seen in Fig. 2, the parameter $\rho_{jk}$ governs the dependence between $\Theta_{j}$ and $\Theta_{k}$. To quantify how the value of $\rho_{jk}$ influences the dependence, we consider the circular correlation coefficient proposed by Rivest1982. This measure is considered here because it is a signed coefficient capable of distinguishing between positive and negative dependence. Let
$\xi_{jk}=\frac{\left|E\left\{e^{i(\Theta_{j}-\Theta_{k})}\right\}\right|-\left|E\left\{e^{i(\Theta_{j}+\Theta_{k})}\right\}\right|}{2},$
then the circular correlation coefficient between $\Theta_{j}$ and $\Theta_{k}$ proposed by Rivest1982 is defined as
$\rho_{jk,R}=\left\{\begin{array}[]{ll}\frac{\xi_{jk}}{\max\left[E\left\{\sin^{2}(\Theta_{j}-\mu_{j})\right\},\ E\left\{\sin^{2}(\Theta_{k}-\mu_{k})\right\}\right]},&E\left\{e^{i(\Theta_{j}-\Theta_{k})}\right\}E\left\{e^{i(\Theta_{j}+\Theta_{k})}\right\}\neq 0,\\ 2\xi_{jk},&E\left\{e^{i(\Theta_{j}-\Theta_{k})}\right\}E\left\{e^{i(\Theta_{j}+\Theta_{k})}\right\}=0.\end{array}\right$.
@@
Figure 3: Plot of the circular correlation coefficients of the TPN distributions, including its copula case, for different values of $\rho_{jk}$, and values of $\kappa_{j},\kappa_{k}\in\{0,0.49,2.45\}$: (a) $\kappa_{j}=\kappa_{k}$; and (b) $\kappa_{j}\neq\kappa_{k}$.
Figure 3 shows the circular correlation coefficients $\rho_{jk,R}$ of the TPN distributions for two cases: $\kappa_{j}=\kappa_{k}$ and $\kappa_{j}\neq\kappa_{k}$, which include the copula case $\kappa_{j}=\kappa_{k}=0$. The figure suggests that the circular correlation coefficient is a strictly increasing function of $\rho_{jk}$. The coefficient spans the entire range (1,1)(-1,1) when $\kappa_{j}=\kappa_{k}$, whereas its range is limited when $\kappa_{j}\neq\kappa_{k}$. This can be inferred from the fact that if the range of $\rho_{jk}$ is extended and $\rho_{jk}=-1$ or $1$ in Definition 1, we have
@@
$\Theta_{k}=\mu_{k}+\operatorname{sign}(\rho_{jk})\,\text{atan}^{*}\left\{\frac{\sin(\Theta_{j}-\mu_{j})}{\cos(\Theta_{j}-\mu_{j})+(\kappa_{k}-\kappa_{j})/R_{j}}\right\}.$
This expression implies that $\Theta_{j}$ and $\Theta_{k}$ have perfect linear relationship when $\kappa_{j}=\kappa_{k}$. When $\kappa_{j}\neq\kappa_{k}$, the relationship between $\Theta_{j}$ and $\Theta_{k}$ depends on $R_{j}$, and the influence of $R_{j}$ increases as $|\kappa_{k}-\kappa_{j}|$ increases. This suggests the advantage of the CTPN distribution, which arises from the copula case $\kappa_{j}=\kappa_{k}=0$, over the TPN distribution in terms of the attainable range of the correlation coefficient. Nonetheless, attention must be paid to the fact that the circular correlation coefficient $\rho_{jk,R}$ reaches 1-1 or $1$ when $\Theta_{j}$ and $\Theta_{k}$ have a linear relationship with probability one; see Rivest1982. Therefore, this coefficient is not designed to measure the strength of dependence between the two variables when they exhibit a nonlinear dependence, which occurs when $\kappa_{j}\neq\kappa_{k}$.
In addition, Fig. 3 shows that the positive and negative signs of $\rho_{jk}$ correspond to positive and negative values of the circular correlation coefficient, respectively. The circular correlation coefficient is zero if and only if $\rho_{jk}=0$.
@@ 5 MCMC Algorithm
5.1 MCMC steps for the TPN
In this section, we explain the algorithm that can be used to obtain posterior samples from the TPN distribution. To deal with the identification problem that the TPN inherited form the PN, we use the approach of Mastrantonio2018, which defines a model for the unidentifiable parameters and a posterior remap is used to obtain the identifiable version of the parameters.
While standard priors can be easily assigned to the parameters $\bm{\mu}$ and $\bm{\kappa}$, care must be taken when specifying a prior for $\bm{\Sigma}$ due to its relationship with $\bm{\Sigma}_{c}$. The Inverse Wishart (IW) distribution is a classical choice for modeling a covariance matrix, but not every valid covariance matrix $\bm{\Sigma}$ yields a valid $\bm{\Sigma}_{c}$. This makes its use problematic unless appropriate adjustments are made. The matrix $\bm{\Sigma}$ is defined over the domain $\mathcal{D}$ of all positive definite matrices, and a subset $\mathcal{D}^{*}$ of this domain can be defined as the set of all positive matrices that, when taking the absolute value of their elements, remain positive definite. We define a new distribution, which we call Truncated IW (TIW), as a distribution supported on $\mathcal{D}^{*}$ whose density is proportional to the IW. The normalization constant of this distribution is not known, but, as we will show, it is not needed in the MCMC algorithm. Write $\bm{\Sigma}\sim TIW(\nu,\bm{\Psi})$ if a random matrix $\bm{\Sigma}$ follows a TIW distribution with degrees of freedom $\nu$ and scale matrix $\bm{\Psi}$.
@@ 5.1 MCMC steps for the TPN
$\bm{\Theta}_{i}$ $=(\Theta_{i,1},\ldots,\Theta_{i,d})^{\top}\stackrel{{\scriptstyle iid}}{{\sim}}\mathrm{TPN}_{d}(\bm{\mu},\bm{\kappa},\bm{\Sigma}),\quad\bm{\Sigma}\sim TIW(\nu,\bm{\Psi})\quad(i=1,\dots,n),$ (20)
$\mu_{j}$ $\stackrel{{\scriptstyle iid}}{{\sim}}CU(0,2\pi),\quad\kappa_{j}\stackrel{{\scriptstyle iid}}{{\sim}}{\cal N}_{1}({m},{v})I_{>0}\quad(j=1,\dots,d),$ (21)
where $CU(0,2\pi)$ is the uniform distribution on the circle $(0,2\pi)$ and ${\cal N}_{1}({m},{v})I_{>0}$ denotes a normal distribution with mean m{m} and variance v{v} truncated to the positive real line. As we will show below, the truncated normal ensures that the full conditional of $\kappa_{j}$ is available in closed form, while for $\mu_{j}$, regardless of the distribution, no closed-form full conditional is available.
To make the inference easier, as is often done when working with the PN, the vectors of latent variables are introduced in the model; define $\bm{R}_{i}=(R_{i,1},\dots,R_{i,d})^{\top}$, $\bm{X}_{c,i}=(X_{c,i,1},\ldots,X_{c,i,d})^{\top}$, and $\bm{X}_{s,i}=(X_{s,i,1},\ldots,X_{s,i,d})^{\top}$ for the $i$-th sample of $\bm{\Theta}_{i}$ as in (6), namely,
$\bm{X}_{c,i,j}$ $=R_{i,j}\cos(\Theta_{i,j}-\mu_{j}),\quad\bm{X}_{s,i,j}=R_{i,j}\sin(\Theta_{i,j}-\mu_{j}).$ (22)
We have that the joint density of $(\bm{R}_{i},\bm{\Theta}_{i})$ can be written as in (7), as
@@
For two of the steps of the MCMC algorithm, we need to use the conditional distribution of $x_{c,i,j}$ given all the other variables in the vector $\bm{x}_{c,i}$, which we indicate as $\bm{x}_{c,i,-j}$, and the conditional distribution of $x_{s,i,j}$ given $\bm{X}_{s,i,-j}=\bm{x}_{s,i,-j}$. Define $\bm{\Lambda}_{c}=\bm{\Sigma}_{c}^{-1}$, $\bm{\Lambda}_{s}=\bm{\Sigma}^{-1}$ and $\bm{\kappa}_{-j}$ as the vector $\bm{\kappa}$ without the $j$-the element. Using standard results of a block matrix inversion and its connection with the conditional mean and variance of a multivariate normal, we know that
$x_{c,i,j}|\bm{x}_{c,i,-j}$ $\sim{\cal N}_{1}(\eta_{c,i,j},\lambda_{c,j}^{-1}),\quad x_{s,i,j}|\bm{x}_{s,i,-j}\sim{\cal N}_{1}(\eta_{s,i,j},\lambda_{s,j}^{-1}),$ (24)
$\eta_{c,i,j}$ $=\kappa_{j}-\lambda_{c,j}^{-1}\bm{\Lambda}_{c,j,-j}\left(\bm{x}_{c,i,-j}-\bm{\kappa}_{-j}\right),\quad\eta_{s,i,j}=-\lambda_{s,j}^{-1}\bm{\Lambda}_{s,j,-j}\bm{x}_{s,i,-j},$ (25)
@@
Latent variable $r_{i,j}$. We can easily see that the full conditional of $r_{i,j}$ is proportional to
$\exp\left\{-\frac{(x_{c,i,j}-\eta_{c,i,j})^{2}}{2\lambda_{c,j}^{-1}}\right\}\exp\left\{-\frac{(x_{s,i,j}-\eta_{s,i,j})^{2}}{2\lambda_{s,j}^{-1}}\right\}r_{i,j}$ (26)
which, by defining $\bm{\eta}_{i,j}=(\eta_{c,i,j},\eta_{s,i,j})^{\top}$, $\tilde{\bm{\Sigma}}_{i,j}=\text{diag}(\lambda_{c,j}^{-1},\lambda_{s,j}^{-1})^{\top}$, and $\bm{x}_{i,j}=(x_{c,i,j},x_{s,i,j})^{\top}$, can be written as
$\exp\left\{-\frac{(\bm{x}_{i,j}-\bm{\eta}_{i,j})^{\top}\tilde{\bm{\Sigma}}_{i,j}^{-1}(\bm{x}_{i,j}-\bm{\eta}_{i,j})}{2}\right\}r_{i,j}$ (27)
which shows that $\bm{\theta}_{i,j}\mid\cdots\sim\mathrm{PN}(\bm{\eta}_{i,j},\tilde{\bm{\Sigma}}_{i,j})$, where $\bm{\theta}_{i,j}\mid\cdots$ indicates the full conditional distribution of $\bm{\theta}_{i,j}$ (i.e., the dots represent the data and all other model components with the exception of $\bm{\theta}_{i,j}$), and hence we can use the algorithm proposed by Hernandez-Stumpfhauser2017. First, we define
$\bm{\xi}_{i,j}=\{\cos(\theta_{i,j}),\sin(\theta_{i,j})\}^{\top},\quad A_{i,j}=\bm{\xi}_{i,j}^{\top}\tilde{\bm{\Sigma}}_{i,j}^{-1}\bm{\xi}_{i,j},\quad B_{i,j}=\bm{\xi}_{i,j}^{\top}\tilde{\bm{\Sigma}}_{i,j}^{-1}\bm{\eta}_{i,j}.$
Then, we draw ${\beta}_{1,i,j}\sim U[0,\exp\{-2^{-1}A_{i,j}(r_{i,j}-B_{i,j}/A_{i,j})^{2}\}]$ and ${\beta}_{2,i,j}\sim U(0,1)$, and the new sample of $r_{i,j}$ is
$r_{i,j}=\left\{(\varrho_{2,i,j}^{2}-\varrho_{1,i,j}^{2}){\beta}_{2,i,j}+\varrho_{1,i,j}^{2}\right\}^{1/2},$
where $U(a,b)$ denotes the uniform distribution on $(a,b)$ and
$\varrho_{1,i,j}=\frac{B_{i,j}}{A_{i,j}}+\max\left\{-\frac{B_{i,j}}{A_{i,j}},-\left(\frac{-2\log{\beta}_{1,i,j}}{A_{i,j}}\right)^{1/2}\right\},\quad\varrho_{2,i,j}=\frac{B_{i,j}}{A_{i,j}}+\left(\frac{-2\log{\beta}_{1,i,j}}{A_{i,j}}\right)^{1/2}.$
Parameter $\mu_{j}$. Being $\mu_{j}$ embedded inside the sine and cosine used to compute $\bm{X}_{c,i}$ and $\bm{X}_{s,i}$, it is not generally possible to find a closed-form expression for its full conditional, which must be then sampled using a Metropolis Hastings algorithm.
Parameter $\kappa_{j}$. Since the kernel of the prior distribution of $\kappa_{j}$ is the same as the one of a normal distribution, from the conditional distribution (24), we can see that
$x_{c,i,j}+\lambda_{c,j}^{-1}\bm{\Lambda}_{c,j,-j}\left(\bm{x}_{c,i,-j}-\bm{\kappa}_{-j}\right)\sim{\cal N}_{1}(\kappa_{j},\lambda_{c,j}^{-1}).$
Hence, the full conditional of $\kappa_{j}$ is proportional to
$\exp\left(-\frac{1}{2}\left[\kappa_{j}^{2}v^{-1}-2\kappa_{j}v^{-1}m+\kappa_{j}^{2}n\lambda_{c,j}-2\kappa_{j}\sum_{i=1}^{n}\left\{\bm{x}_{c,i}+\lambda_{c,j}^{-1}\bm{\Lambda}_{c,j,-j}\left(\bm{x}_{c,i,-j}-\bm{\kappa}_{-j}\right)\right\}\right]\right)$ (28)
$\propto\exp\left\{-\frac{(\kappa_{j}-m_{p,j})^{2}}{2v_{p,j}}\right\},$ (29)
@@
$v_{p,j}$ $=(v^{-1}+n\lambda_{c,j})^{-1},\quad m_{p,j}=v_{p,j}\left[\sum_{i=1}^{n}\left\{\bm{x}_{c,i}+\lambda_{c,j}^{-1}\bm{\Lambda}_{c,j,-j}(\bm{x}_{c,i,-j}-\bm{\kappa}_{-j})\right\}+v^{-1}m\right].$
It is easy to see that this is the kernel of ${\cal N}_{1}(m_{p,j},v_{p,j})$, and since $\kappa_{j}>0$, the full conditional of $\kappa_{j}$ is ${\cal N}_{1}(m_{p,j},v_{p,j})\bm{I}_{\kappa_{j}>0}$.
Parameter $\bm{\Sigma}$. To sample this parameter, we use a Metropolis–Hastings step. At iteration $b$, the proposal distribution is set to $IW(\tau,\bm{\Sigma}^{b}(\tau-(d+1)))$, where $\bm{\Sigma}^{b}$ denotes the most recently accepted value of the parameter in the MCMC and $IW(\tau,\bm{\Sigma}^{b}(\tau-(d+1)))$ indicate the Inverse Wishart distribution with degrees of freedom $\tau$ and scale matrix $\bm{\Sigma}^{b}\{\tau-(d+1)\}$. The expected value of this proposal distribution is $\bm{\Sigma}^{b}$, and as the parameter $\tau$ increases, the variance of the distribution decreases, making $\tau$ a tuning parameter that controls the variability of the proposal.